Prepare journal entries to record the following selected transactions that occurred during the company's operations: (4 marks)

1.     On January 30, 2017, Candlestick, Inc. sells $100,000, five-year, 10 % bonds for 92.639% of face value.

2.     Company sells $100,000, five-year, 10 % bonds for 108.111% of face value on July 1, 2017. 

1
Expert's answer
2021-07-12T11:45:19-0400

1.). Determine the cash received on bonds issuance = 100,000 x 92.639"\%" = 92,639

Discount on bonds payable = 100,000 – 92,639 = 7,361

Determine Bond interest expense:

Bond interest = 100,000 "\times" 10"\%" = 10,000"\times"5 years = 50,000

Total bonds value plus interest over 5 years = 100,000 + 50,000 = 150,000

Bonds interest expense = 150,000 – 92,639 = 57361


Amortized bonds interest expense for 5 years = "\frac{57,361}{5}" = 11,472.20

Discount amortization = 11,472.20 - 10,000 = 1,472.20


Journal entries:

Dr. Cash                                       92,639

Dr. Discount on bonds payable     7,361   

               Cr. Bonds payable                                100,000

(To record issue of 10"\%", 5-year bonds at a discount)       

 

Dr. Bonds interest expense                         11,472.20                          

               Cr. Discount on bonds payable                             1,472.20   

               Cr. Cash                                                                 10,000

(To record periodic interest payments)     

 

Dr. Bonds payable                 100,000

              Cr. Cash                                              100,000

(To record payment of bonds face value at maturity)    


2.). Determine the cash received on bonds issuance = 100,000"\times"108.111"\%" = 108,111

Premium on bonds payable = 108,111 – 100,000 = 8,111

Determine Bond interest expense:

Bond interest = 100,000 "\times"10"\%" = 10,000"\times"5 years = 50,000

Total bonds value plus interest over 5 years = 100,000 + 50,000 = 150,000

Bonds interest expense = 150,000 – 108,111 = 41,889


Amortized bonds interest expense for 5 years = "\frac{41,889}{5} = 8,377.80"


Premium amortization = 10,000 – 8,377.80 = 1,622.20


Journal entries:

Dr. Cash                                                     108,111

               Cr. Premium on bonds payable                                8,111   

               Cr. Bonds payable                                                 100,000

(To record issue of 10%, 5-year bonds at a premium)       

 

Dr. Bonds interest expense                            8,377.80                          

Dr. Discount on bonds payable                     1,622.20   

                   Cr. Cash                                                                10,000

(To record periodic interest payments)     

 

Dr. Bonds payable                                          100,000

              Cr. Cash                                                                   100,000

(To record payment of bonds face value at maturity)